INFRARED SPECTRUM OF H_{3}^{+} NEAR THE GALACTIC CENTER

\underline{TAKESHI OKA}, Department of Astronomy and Astrophysics, and Department of Chemistry, The Enrico Fermi Institute, The University of Chicago, Chicago IL 60637.

The central 150 pc of the Galaxy called the Central Molecular Zone (CMZ) is the hub of astrophysical activities. It has high densities of stars, interstellar matter, supernovae, gravitational and magnetic energies, and emit intense radio and X-rays. A large fraction (∼ 10 %) of molecules in the Galaxy are concentrated in this region. Nevertheless its extinction is small since molecules are concentrated in dense (≥ 10^{4} cm^{−3}) and massive molecular clouds with a volume filling factor of only ∼ 0.01. What kind of gas fills the rest of the CMZ has been a vexed question. Our observations using infrared spectrum of H_{3}^{+} have revealed a vast amount of low density (∼ 100 cm^{−3}) and high temperature (∼ 250 K) gas with a high volume filling factor.^{a} Crucial to this discovery has been the unique characteristics of H_{3}^{+} as a probe based on its simple chemistry and spectroscopy. We use it as a dosimeter, a depth meter, a thermometer, and a densitometer to measure ionization rate ζ, dimensions of the gas L, temperature T, and density n, respectively.^{b} The H_{3}^{+} spectrum in the CMZ first appeared as an enigmatic, intense and broad line toward the brightest star. The line was more than 10 times stronger than any H_{3}^{+} line in the Galactic disk indicating a large amount of H_{3}^{+} but we did not know how to interpret their intricate velocity profile. Our discovery in 2002 of the spectrum of H_{3}^{+} in the J = K = 3 metastable rotational level^{c} provided the key, which we used as the Rosetta stone to decipher the enigmatic spectrum. Now observations of H_{3}^{+} toward more than dozen stars from the center to 30 pc to the East and their analyses based on a model calculation of the thermalization of H_{3}^{+}^{d}, have established existence of a large amount of the warm and diffuse gas. The revelation of this new category of gas drastically changes the previous concept of the gas in the CMZ.^{e} Its relation with the previously postulated three categories of gas will be discussed.

aT. Oka, T. R. Geballe, M. Goto, T. Usuda, & B. J. McCall, ApJ, 632, 882 (2005)

bT. Oka, Proc. Natl. Acad. Sci. USA, 103, 12235 (2006)

cM. Goto, B. J. McCall, T. R. Geballe, T. Usuda, N. Kobayashi, H. Terada, & T. Oka, PASJ, 54, 951 (2002)

dT. Oka & E. Epp, ApJ, 613, 349 (2004)

eGoto, Usuda, Nagata, Geballe, McCall, Indriolo, Suto, Henning, Morong, & Oka, ApJ submitted.